OpenStudy (anonymous):

How to simplify (x+3 / x-2) + (x+1 / x-1) = 10/3 so I can insert values for -b +/- sqrt (b^2 - 4ac / 2a) a quadratic formula

OpenStudy (***[isuru]***):

hi, \[\frac{ x +3 }{x -2 } + \frac{ x + 1 }{ x -1 } = \frac{ 10 }{ 3 }\] multiply each term by ( x-2) \[[(\frac{ x +3 }{ x -2 }) \times ( x-2) ] + [ (\frac{ x +1 }{ x-1}) \times ( x-2) ] = \frac{ 10(x-2) }{ 3}\] now.. multiply each term by ( x-1) \[(x + 3)(x -1) + [ (\frac{ (x +1)(x -2) }{ x-1 }) \times ( x-1)] = \frac{ 10(x-2)(x-1) }{ 3 }\] now... multiply each term by 3 \[3(x + 3)(x-1) + 3(x +1)(x -2) = 10( x-2 )(x -1)\] now.. simplify \[3(x^{2} +2x -3) + 3( x^{2} -x -2 ) = 10(x^{2} -3x +2)\] \[(3x^{2} +6x -9) + ( 3x^{2} -3x -6 ) = (10x^{2} -30x +20)\] \[3x^{2} +6x -9+ 3x^{2} -3x -6 = 10x^{2} -30x +20\]\[10x^{2} - 3x^{2} - 3x^{2} -30x -6x +3x +20 +9 +6 = 0\]\[4x^{2} -33x +35 = 0\] now u can apply the formula.. hope this will help ya!!!

OpenStudy (***[isuru]***):

did ya got it ? :)

OpenStudy (anonymous):

I get it now, but what will a = b= and c= be?

OpenStudy (***[isuru]***):

a = 4 b = -33 c = 35 but before applying the values , solve the problem by urself and check whether the values r same... :P ....